OpenStudy (anonymous):

Write in Logrithmic exponents log_a^5=7

OpenStudy (anonymous):

is it \[\log_{10} \]?

sam (.sam.):

\[\huge \log _{a}5=7~~~~???\]

OpenStudy (anonymous):

Yes I have to write in logrithmic expressions.

OpenStudy (anonymous):

it is already in logarithmic form maybe you have to write it in exponential form?

OpenStudy (anonymous):

\[\log_b(x)=y\iff b^y=x\] so you your case you would have \[\log_a(5)=7\iff a^7=5\]

Directrix (directrix):

Mathematics has its own language and learning to use it is not always easy. Just anything will not do. A dash becomes a negative sign. A set of dots .... could mean something more than just dots. I'm not sure how to interpret your given problem but I uploaded a reminder on the definition of logarithm that may help.